# RODS: Robust Optimization Inspired Diffusion Sampling for Detecting and Reducing Hallucination in Generative Models

### v1

@@ A.1 Continuation Method and sampling process matching
Thus, at $t=T$, minimizing $f_{T}(x)$ aligns with minimizing $-\log p_{T}(x)$, which confirms the behavior of the diffusion process in this case.
@@
Therefore, minimizing $f_{t}(x)$ over $x$ is equivalent to minimizing $-\log p_{t}(x)$ over $x$.
@@ A.2 Understanding of the index
To intuitively understand when the learned score function becomes unstable, we look at how rapidly it changes in space. Let $v(x):=\nabla_{x}\log P_{t}(x)$ be the score function at time $t$, pointing in the direction of increasing data likelihood. In well-behaved regions, this vector field changes smoothly. But in low-density areas, $v(x)$ can shift direction quickly—these are the regions where hallucinations are likely to occur.
@@ B.1 Dataset and Base Model
We evaluate our method on three benchmarks: AFHQ-v2 [23], FFHQ [23], and 11k-hands [1]. AFHQ-v2 contains high-resolution animal face images across three domains (cats, dogs, wild), with improved alignment and quality compared to the original AFHQ. FFHQ includes 70,000 high-quality human face images with rich variation in age, ethnicity, and lighting, making it a standard benchmark for face synthesis. The 11k-hands dataset contains 11,076 normal hand images (1600$\times$1200 pixels) from 190 individuals aged 18–75. Each subject was photographed opening both hands, from dorsal and palmar views, against a white background with consistent distance from the camera.
@@ B.2 Detection Threshold
Figure 6: ROC curve for hallucination detection on the FFHQ dataset. Each point corresponds to a different detection threshold. Lower thresholds yield higher true positive rates but also increase false positives. For example, a threshold of 0.09 achieves $\sim$80TPR at under 20% FPR. Thresholds can be adjusted depending on the application’s sensitivity to hallucination risk.
@@ B.5 Acceleration Strategy and Running Statistics
Method Hall.$\downarrow$ Correction%$\uparrow$ New Hall.$\downarrow$ Time (s)
@@ B.6 Labeling Detail
Since there is no universally agreed-upon definition of hallucination in generative models, and no established gold standard for evaluation, we resort to manual annotation to assess the presence of hallucinations. Given that the identification of visual hallucinations in our setting does not require domain-specific expertise, we enlisted lab members to perform the labeling.
@@
Examples of annotated hallucinations can be found in Figure 11 (faces) and Figure 12 (hands).
@@ C.3 Reduce Hallucination in Diffusion Model
Trajectory variance filtering. Aithal et al. [2] observe that hallucinated samples often exhibit high prediction variance in the reverse diffusion trajectory, signaling off-manifold behavior. They propose a simple yet effective variance-based metric to filter out such aberrant samples.

### v2

@@ A.1 Continuation Method and sampling process matching
Thus, at $t=T$, minimizing $f_{T}(x)$ aligns with minimizing $\log p_{T}(x)$, which confirms the behavior of the diffusion process in this case.
@@
Therefore, minimizing $f_{t}(x)$ over $x$ is equivalent to minimizing $\log p_{t}(x)$ over $x$.
@@ A.2 Understanding of the index
To intuitively understand when the learned score function becomes unstable, we look at how rapidly it changes in space. Let $v(x):=\nabla_{x}\log P_{t}(x)$ be the score function at time $t$, pointing in the direction of increasing data likelihood. In well-behaved regions, this score field changes smoothly. But in low-density areas, $v(x)$ can shift direction quickly—these are the regions where hallucinations are likely to occur.
@@ B.1 Dataset and Base Model
[…]
@@ B.2 Detection Threshold
In our current experiments, we observed that the choice of $\epsilon$ is influenced by both the dataset characteristics and the desired sensitivity level. For instance, 11k-hands features clean backgrounds and fine-grained hand details, so a smaller $\epsilon$ helps detect subtle hallucinations in localized regions. In contrast, face datasets (AFHQv2, FFHQ) often involve more complex backgrounds, where a slightly larger $\epsilon$ is preferred to avoid overreacting to natural variation in curvature.
@@
Figure 6: ROC curve for hallucination detection on the FFHQ dataset. Each point corresponds to a different detection threshold. Lower thresholds yield higher true positive rates but also increase false positives. For example, a threshold of 0.09 achieves $\sim$80% TPR at under 20% FPR. Thresholds can be adjusted depending on the application’s sensitivity to hallucination risk.
@@ B.5 Acceleration Strategy and Running Statistics
This result also explains why simply reducing the step size is insufficient. Even with 200 uniform steps, which is five times smaller than the 40-step baseline, the hallucinations persist. In contrast, RODS-CAS adaptively adjusts its update direction in high-curvature regions and eventually removes artifacts. The key issue lies in directional error within low-density, high-uncertainty areas; smaller steps cannot fix a wrong direction, while curvature information helps steer sampling toward stable regions.
@@
[…]
@@ B.6 Labeling Detail
Since there is no universally agreed-upon definition of hallucination in generative models, and no established gold standard for evaluation, we resort to manual annotation to assess the presence of hallucinations. Given that the identification of visual hallucinations in our setting does not require domain-specific expertise, we recruited lab members to perform the labeling.
@@
To further improve evaluation precision, we conducted pairwise human comparisons on anonymized image pairs (RODS vs. baseline). Unlike binary hallucination labels, these pairwise judgments capture relative image quality, providing a more nuanced measure of generation fidelity. The evaluation rubric is as follows:
  • Better: The RODS-generated image is perceived as more faithful or realistic, often due to improved correction of visual anomalies (e.g., fixing a misplaced eye) without introducing new defects.
  • Worse: The RODS-generated image is clearly degraded relative to the baseline, typically because it introduces new artifacts or hallucinated elements that were not present in the original.
  • Same: No human-identifiable difference exists between the baseline and the RODS-generated image.
  • Unclear: A change occurs, but it does not involve new hallucinations, and it is difficult to judge which version is better (e.g., slight pose changes or missing accessories like a ring).
All images were labeled independently by two human raters to ensure objectivity. Whenever their judgments diverged, the pair reviewed the cases together and discussed until they reached a consensus, producing a single, agreed-upon label for each image. This consensus process both minimized individual bias and improved the overall reliability of our ground-truth data.To evaluate RODS-CAS’s impact relative to the baseline (EDM-Euler), annotators first identified hallucinations in the Euler-sampled outputs. They then compared each corrected image against its Euler counterpart, categorizing the result as “better,” “worse,” “same," or "unclear". This side-by-side comparison made it easy to see whether our method truly reduced artifacts, or introduced new ones. Examples of annotated hallucinations can be found in Figure 11 (faces) and Figure 12 (hands).
@@ C.3 Reduce Hallucination in Diffusion Model
Trajectory-based consistency control. Aithal et al. [2] observe that hallucinated samples often exhibit high prediction variance in the reverse diffusion trajectory, signaling off-manifold behavior. They propose a simple but effective variance-based metric to filter out such aberrant samples. Huang et al. [22] regularizes the sampling trajectory by constraining the norm of the predicted noise to remain aligned with the score direction, ensuring update validity within high-confidence regions.
@@
Attention Modulation Adjustment. Oorloff et al.  [38] propose Adaptive Attention Modulation (AAM) to reduce hallucinations in diffusion models by dynamically adjusting self-attention behavior during inference. They apply masked perturbations to disrupt early-stage noise that may propagate into hallucinations. AAM directly modulates the internal attention layers, providing a lightweight way to suppress hallucinated artifacts while preserving image quality.